A 0.105-kg hockey puck moving at 30m/s is caught and held by a 79-kggoalie at rest. With what speed does the goalie slide on theice?
m/s
Answer #1
Applying the law of conservation of momentum we have as,
m1v1=m2v2
0.105*30=79*v
v=3.15/79m/s
Hence we get by it.